OpenStudy (anonymous):

HELP ME MY HEAD IS ABOUT TO BLOWUP q q - / - 7 26

OpenStudy (anonymous):

make your head blow up please lol

OpenStudy (anonymous):

HAHA funny im seriose plz i will give madle

OpenStudy (anonymous):

yes. lol

OpenStudy (lyrae):

\[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

that doesnt help plz im in the middle of an exam and its timed so i have no time

OpenStudy (anonymous):

um then close this. it is a test www.openstudy.com/code-of-conduct/

OpenStudy (anonymous):

copy and paste. but first close this

OpenStudy (anonymous):

no its pactice exam

OpenStudy (anonymous):

ya though you worded it different. be careful. they can suspend u for tht

OpenStudy (anonymous):

ok thx 4 letting me know

OpenStudy (lyrae):

If you replace a, b, c, d with your own values it should be quite easy to simplify.

OpenStudy (anonymous):

is it q^2/182

OpenStudy (lyrae):

\[\frac{ \frac{ q }{ 7 } }{ \frac{ q }{ 26 } } = \frac{ q }{ 7 } \times \frac{ 26 }{ q }\]

OpenStudy (kohai):

And then from there you cross multiply and then square the number to find q

OpenStudy (anonymous):

thx

OpenStudy (lyrae):

Yw :)